OpenStudy (anonymous):

1. Let f(x) = x2 – 81. Find f–1(x). 2. Let f(x) = x + 2 and g(x) = x2 – 6x + 3. Find g(f(x)). 3. Let f(x) = x – 3 and g(x) = x + 11. Find f(x) • g(x). 4. Let f(x) = x2 – 3x – 7. Find f(–3). 5. Let f(x) = x2 + 2x – 5 and g(x) = x + 6. Find f(x) + g(x). 6. Let f(x) = x2 – 2x + 2 and g(x) = x – 3. Find f(x) • g(x). 7. Let f(x) = –4x + 7 and g(x) = 10x – 6, find f(g(x)). 8. Let f(x) = 5x + 12. Find f–1(x). 9. Let f(x) = x – 2 and g(x) = x2 – 7x – 9. Find f(g(–1)). 10. Let f(x) = 3x2 + x – 3 and g(x) = x2 – 5x + 1. Find f(x) – g(x).

OpenStudy (neo92):

are u sure that da questions (1) & (8) are correct? anyway (2) just f(x) for x in g(x) i.e g(f(x)) = (x+2)^2 - 6(x+2) +3 question (7) also can be solved in the same way. (3). f(x).g(x) = f(x)*gx) = (x-3)(x+11) = x^2+8x-33 (4). just substitute (-3) for x (5). just add the two functions (6). same procedure as in (3) (7). same as in (2) (9). first substitute (-1) for x in g(x) and find g(-1) then substitute the value obtained for g(-1) for x in f(x) (10).subtract g(x) from f(x)

OpenStudy (anonymous):

Thanks!